OpenStudy (anonymous):

What is the lim (1-cos theta)/(2sin^2 theta) as theta approahes 0?

OpenStudy (anonymous):

Do you know the result \[\lim_{x \rightarrow 0}sinx / x =1\]

OpenStudy (anonymous):

yeah..thts an identity

OpenStudy (anonymous):

\[\lim_{\Theta \rightarrow 0} (1-\cos \Theta) / (2\sin ^{2}\Theta)\]

OpenStudy (anonymous):

So using this result you get ehe answer, \[1-\cos \theta=2 \sin ^{2}(\theta/2)\]

OpenStudy (anonymous):

wait y are they equal

OpenStudy (anonymous):

That an identity do you know cos 2theta=cos ^2 theta - sin^2 theta?

OpenStudy (anonymous):

oh ok..srry i forgot

OpenStudy (anonymous):

Ok yeah so you have sin^2 (x/2) / sin^ x (x is more easier to type than theta) so multiply and divide by x^2 , so you have 4 sin^2 (x/2)/(x/2)^2/ (sin^2 x/x^2) now everything except 4 becomes 1. The answer is 4. Can you tell me why?

OpenStudy (anonymous):

wen i do it in teh calculator it says the answer is 1/4....cept im not sure why

OpenStudy (anonymous):

Oh yeaah i made a mistake you have (x/2)^2 in the denominator right so you have to divide by 4 and not mulyipli i'm sorry the answer is 1/4. Its like pretty hard to explain over computer just write it down. You'll understand

OpenStudy (anonymous):

If you have sin (x/2) / x as x tends to 0 you have to first make x x/2 so you get 1/2 in the denominator similarly try to solve the given problem like this only there is an x^2 so you have to mutiply and divide by 4 you will get 1/4 only.

OpenStudy (anonymous):

can u explain it the way wolframalpha did it..i understand evrything in it ecxpet wen it applied L Hopitals

OpenStudy (anonymous):

im jsut askin u to explauiin it taht way b.c its already typed mathematically on teh comp

OpenStudy (anonymous):

Do you know L hospitals rule its just another way of solving the problem do you know L hospitals rule?

OpenStudy (anonymous):

no

OpenStudy (anonymous):

Ok see if you have any function such as f(x)/g(x) and both f(x) and g(x) tend to 0 then in that case you can just get the limit by differentaiting numerator and denominator eg., take sin x /x as x ->0 both numerator and denominator tend to 0 so differentiate numerator and denominator you get cos x /1. Now at x=0 cos x=1 so the net limit is 1.

OpenStudy (anonymous):

Basically \[\lim_{f(x) \rightarrow 0\lim_{g(x) \rightarrow 0}} f(x)/g(x)= f'(x)/g'(x)\]

OpenStudy (anonymous):

Did you understand so far?

OpenStudy (anonymous):

yes

OpenStudy (anonymous):

ok i get it....u diffrenitate to get it so teh limit is not 0/0

OpenStudy (anonymous):

So now in our question 1-cos x tends to 0 as x->0 since cosx=1 and denominator also tends to 0 as sin^2 x tends to 0 as sin x=0, so differentiate both numerator and denominator

OpenStudy (anonymous):

So now apply L hospitals rule and tell me what you get....

OpenStudy (anonymous):

1/ 2cos (x)

OpenStudy (anonymous):

no you get 1/4cos x check again there was already a 2 in the denomintor of the question.

OpenStudy (anonymous):

oh yeah 1/4 (lim x->0 of 1/cos x)

OpenStudy (anonymous):

thats 1 as cos x=1

OpenStudy (anonymous):

yeah u get 1/4 sicne as cos x goes to 0 u get 1

OpenStudy (anonymous):

Thanls for ur help! =)

OpenStudy (anonymous):

No problem...